# ToMAP: Training Opponent-Aware LLM Persuaders with Theory of Mind

### v1

@@ 1 Introduction
Persuasion—the process of influencing others’ attitudes, beliefs, or behaviors—is fundamental to human communication and social interaction, and is recognized as a complex and advanced social behavior due to its high demands on reasoning and cognition. Throughout history, countless orators and scholars have explored the art of persuasion, seeking techniques to effectively shape others’ opinions, a pursuit that large language models (LLMs) are now beginning to undertake as well. State-of-the-art LLMs like GPT-4o have demonstrated remarkable persuasive abilities, generating persuasive arguments [1, 2] and successfully influencing human attitudes in specific domains [3, 4].
Despite these advances, current LLM-based persuaders face significant limitations in modeling the interactive and dynamic nature of persuasion. During a conversation, human participants naturally consider related claims around the central claim, forming a rich cognitive graph as shown in Figure 1. Intuitively, skilled human persuaders navigate this cognitive graph, analyzing the persuadee’s mental state and identifying “nodes” in the cognitive graph that are ripe for elaboration and could potentially affect the central claim. This process relies on the ability to model and reason about others’ beliefs, a cognitive capacity known as theory of mind (ToM). As a cornerstone of human social cognition, ToM enables humans to tailor their messages to the opponent’s perspectives dynamically, making persuasion more effective and opponent-aware.
@@
LLMs, by contrast, often overlook the interconnected structure of the cognitive graph and rarely consider the opponent’s perspective in dynamic exchanges. This deficiency in theory of mind leads to two key shortcomings in current LLM-based persuaders. First, after stating their position in the initial round, they tend to repeat the same arguments rather than drawing on new insights from the broader cognitive graph [5]. For instance, the persuader trained without ToM modules exhibits up to 12% content overlap between consecutive turns (Footnote 7). Second, LLM-based persuaders are often rigid and self-centered, focusing narrowly on their own arguments rather than adapting to the opponent’s views dynamically [6, 7]. This lack of flexibility is exemplified in Figure 6, where baseline models employ opponent-oriented strategies far less frequently.
@@
Our experiments demonstrate that the ToMAP model, with only 3B parameters, outperforms baseline configurations and exhibits persuasiveness comparable to or exceeding that of several state-of-the-art LLMs, including GPT-4o (39.4% relative gain) and Deepseek-R1 (2.7% relative gain) across diverse corpora and against multiple persuadee models. Ablation studies underscore the importance of ToM modeling and the RL process, as ablating either leads to significant drops in the persuasion effect. Further analyses reveal that ToMAP produces more complex reasoning chains and reduces repetition during training, achieves steady persuasion gains in long conversations, and adopts more logical and opponent-aware strategies. These findings highlight ToMAP’s unique strengths robust and effective framework for developing more persuasive LLM agents, suggesting theory of mind and opponent modeling as promising directions for building more persuasive language agents.
@@ 2 Related Work
LLM Persuasion. Persuasion, broadly defined as influencing others’ attitudes, beliefs, or behaviors through communication [8], is a fundamental ability of humans. Recently, LLM persuasion has attracted increasing attention with the rapid development of state-of-the-art LLMs [9, 10, 11, 12]. Researchers found LLMs can produce persuasive content similar to, or even surpassing, human experts on certain areas such as political issues [1, 2, 13], vaccine promotion [3], investment decisions [4], and factual knowledge [14]. Moreover, empirical evidence shows that LLMs can potentially shift human attitudes in real-world conversations [15, 16, 13]. More recently, several systematic frameworks have been proposed to evaluate the persuasiveness of LLMs, either by comparing two arguments [17] or measuring the attitude and behavioral change of the persuadee model in conversations [15, 18, 19, 20].
A closely related line of research focuses on enhancing the persuasion capabilities of LLMs. [21, 22] summarize common human persuasion techniques, which subsequent studies leverage to train LLMs explicitly in these strategies [23]; [24] explores retrieving external evidence to support persuasive arguments, while [19, 25] develop LLMs using high-quality persuasion dialogues. [26]’s work shows model size yields diminishing returns for persuasiveness, indicating simply scaling model size may not make better persuaders. Reinforcement learning (RL) is another common approach to enhance a model’s persuasiveness without annotated data. This line of research has primarily focused on narrow tasks, such as negotiation in games [27, 28] or specific traits like consistency [29]. Concurrently, [30] explores training general persuasion models with RL and utilizes a Bayesian persuasion framework to evaluate persuasiveness.
Theory of Mind. Theory of Mind (ToM) is a crucial concept in psychology, which refers to the ability to understand others by attributing mental states such as beliefs, desires, intentions, and emotions [31, 32, 33, 34]. ToM enables individuals to predict and interpret others’ behaviors in society and facilitates complex communication, making it a cornerstone of human meta-cognition ability. ToM ability in AI systems has attracted great attention [35, 36, 37, 38], and researchers have proposed benchmarks to evaluate LLMs’ ability to reason over others’ mental state [39, 40, 41, 42, 43]. More complex ToM tasks, namely ToM on fiction scenarios [44] and high-order ToM [45, 46], are also introduced, showing future directions of enhancing LLMs’ ToM abilities.
The relationship between persuasion and Theory of Mind (ToM) has been widely studied [47, 48, 49], as understanding another’s mental state is fundamental to influencing it. More recently, [50] evaluated LLMs’ ToM in persuasive dialog, and [51] highlighted the lack of ToM in LLM conversations. Our work also explores the intersection of persuasion and ToM, training LLMs to utilize ToM information more effectively to enhance persuasion techniques.
@@ 3 Methodology
In this section, we introduce the key components of Theory of Mind Augmented Persuader (ToMAP). We first describe the setting of the persuasion task in Section 3.1 and the reinforcement learning algorithm used to train the persuader in Section 3.2. Finally, we introduce two theory of mind modules that provide critical information about the opponent’s thoughts Section 3.3.
@@ 3.1 The Setting of Persuasion Task
The persuasion process is a multi-turn conversation between two LLM agents: a persuader, denoted as $\mathbf{X}$, and a persuadee, denoted as $\mathbf{Y}$. During the conversation, the persuader aims to convince the persuadee to support a given claim $Q$. Formally, we denote the $i$th turn in the conversation as $T_{i}$, and the conversation history till the $i$th turn as $H_{i}=T_{1\ldots i}$ [^1] Specially, we define $H_{0}$ as an empty string.. Since the two agents take turns speaking in the conversation, the next turn can be represented as
$T_{n}\sim\begin{cases}P_{\mathbf{X}}(\cdot\mid p,Q,H_{n-1})&\quad\text{if }n\text{ is odd,}\\ P_{\mathbf{Y}}(\cdot\mid p,Q,H_{n-1})&\quad\text{if }n\text{ is even,}\end{cases}$ (1)
where $p$ is the task-specific prompt (refer to Appendix B).
After each persuadee’s turn, we employ a 5-point Likert scale [52] to assess the persuadee’s perspectives [30, 18], represented as:
$s^{\prime}(\mathbf{Y},H_{n},Q)\in\{0,1,2,3,4\}.$ (2)
As LLMs are known to exhibit inconsistency in opinions, which means a LLM might agree to both one claim and the opposite claim [53, 54], we consider the persuadee’s opinion on two contradictory claims and calculate the balanced agreement score as:
@@
where $\neg$ means the opposition of the original claim. From Equation 3, we can derive $S(\mathbf{Y},H_{n},Q)+S(\mathbf{Y},H_{n},\neg Q)=1$, which ensures the consistency and reliability of LLM opinion judgment.
@@ 3.2 A Reinforcement Learning Formulation for Persuader Training
Reinforcement learning (RL) has demonstrated great success in recent work [55, 56, 57], where models develop complex reasoning skills. Similar to  [30], we utilize RL to train the persuader through to incentivize LLMs’ persuasion ability without extensive labeled data.
As we find directly optimizing the persuader in multi-round interaction is challenging and requires much more complex design [56, 58], we propose decomposing the multi-turn conversations as single-turn activities to enhance the stability and efficiency of our training schema. Specifically, given the conversation history $H$ and target claim $Q$, we optimize the persuader to generate the next argument that maximizes the attitude shift of the persuadee. Formally, at the persuader’s turn $12k-$ 1, we use
@@
to measure the attitude shift after the persuader’s turn and the persuadee’s following turn. Then, we normalize the agreement gap to the range of $[-1,1]$ to obtain the persuasion reward:

### v2

@@ 1 Introduction
Persuasion—the process of influencing others’ attitudes, beliefs, or behaviors—is fundamental to human communication and social interaction, and is recognized as a complex and advanced social behavior. Recently, large language models (LLMs) have also begun to engage in this cognitively demanding practice. State-of-the-art LLMs like GPT-4o have demonstrated remarkable persuasive abilities, which can generate persuasive arguments (Bai et al., 2023; Palmer & Spirling, 2023) and influence human attitudes in specific domains (Karinshak et al., 2023; Takayanagi et al., 2025). At the same time, as LLMs become increasingly integrated into diverse social and professional domains like customer servicing and proofreading, effectively persuading an LLM has also grown in importance.
Despite the advances in LLM persuasion, current LLM-based persuaders face significant limitations in modeling the interactive and dynamic nature of persuasion. During a conversation, human participants naturally consider related claims around the central claim, forming a rich cognitive graph as shown in Figure˜1. Intuitively, skilled human persuaders navigate this cognitive graph, analyzing the persuadee’s mental state and identifying “nodes” in the cognitive graph that are ripe for elaboration and could potentially affect the central claim. This process relies on the ability to model and reason about others’ beliefs, a cognitive capacity known as theory of mind (ToM). As a cornerstone of human social cognition, ToM enables humans to tailor their messages to the opponent’s perspectives dynamically, making persuasion more effective and tailored to their opponents.
@@
LLMs, by contrast, often overlook the interconnected structure of the cognitive graph and rarely consider the opponent’s perspective in dynamic exchanges. This deficiency in theory of mind leads to two key shortcomings in current LLM-based persuaders. First, after stating their position in the initial round, they tend to repeat the same arguments rather than drawing on new insights from the broader cognitive graph (Chu et al., 2024). For instance, the persuader trained without ToM modules exhibits up to 12% content overlap between consecutive turns (Footnote˜8). Second, LLM-based persuaders are often rigid and self-centered, focusing narrowly on their own arguments rather than adapting to the opponent’s views dynamically (Wang et al., 2025a; Becker, 2024). This lack of flexibility is exemplified in Figure˜6, where baseline models employ opponent-oriented strategies far less frequently.
@@
Our experiments demonstrate that the ToMAP model, with only 3B parameters, significantly outperforms baseline configurations and exhibits persuasiveness comparable to or exceeding that of several state-of-the-art LLMs across diverse corpora and against multiple persuadee models. Further analyses reveal that ToMAP produces more complex reasoning chains and reduces repetition during training, achieves steady persuasion gains in long conversations, and adopts more logical and opponent-aware strategies. These findings highlight ToMAP’s unique strengths robust and effective framework for developing more persuasive LLM agents, suggesting theory of mind and opponent modeling as promising directions for building more persuasive language agents.
@@ 2 Related Work
LLM Persuasion. LLM persuasion has attracted increasing attention with the rapid development of state-of-the-art LLMs (Bozdag et al., 2025b; Jaech et al., 2024; Rogiers et al., 2024; Breum et al., 2024), showing great potentials on areas like political issues (Bai et al., 2023; Palmer & Spirling, 2023; Potter et al., 2024), vaccine promotion (Karinshak et al., 2023), investment decisions Takayanagi et al. (2025), and factual knowledge (Costello et al., 2024). Empirical evidence also shows that LLMs can potentially shift human attitudes in real-world conversations (Salvi et al., 2024; Shi et al., 2020b; Potter et al., 2024). Sveral systematic frameworks have been proposed to systematically evaluate the persuasiveness of LLMs (Pauli et al., 2024; Salvi et al., 2024; Bozdag et al., 2025a; Singh et al., 2024; Durmus et al., 2024). Carrasco-Farre (2024) further found the persuasiveness of LLMs comes from proposing arguments that require higher cognitive effort.
To enhance the persuasion capabilities of LLMs, researchers utilize human’s strategies(Wang et al., 2019; Yang et al., 2019) and high-quality dialogues(Singh et al., 2024; Stengel-Eskin et al., 2024; Jin et al., 2024; Furumai et al., 2024) to train LLMs. Reinforcement learning (RL) for persuasion has primarily focused on narrow tasks, such as negotiation games (Keizer et al., 2017; Lewis et al., 2017) or consistency checking (Shi et al., 2020a). Concurrent to our work, Anonymous (2025) explores training general persuasion models with RL.
Theory of Mind. Theory of Mind (ToM) refers to the ability to understand others by attributing mental states (Wimmer & Perner, 1983; Baron-Cohen et al., 1985; Frith & Frith, 2005; Wellman, 2018). ToM ability in AI systems has attracted great attention Cuzzolin et al. (2020); Langley et al. (2022); Williams et al. (2022); Schossau & Hintze (2023), and researchers have proposed benchmarks to evaluate such capabilities (Kim et al., 2023; Le et al., 2019; Wu et al., 2023). More complex ToM tasks are also introduced, showing future directions of enhancing LLMs’ ToM abilities (Xu et al., 2024; He et al., 2023; Street et al., 2024).
The relationship between persuasion and Theory of Mind has been widely studied (Peterson et al., 2018; McAlister & Cornwell, 2009; Slaughter et al., 2013), as understanding another’s mental state is fundamental to influencing it. More recently, Yu et al. (2025) evaluated LLMs’ ToM in persuasive dialog, and Zhang & Zhou (2025) identified the lack of ToM in LLM conversations, highlighting the importance of ToM-enhanced persuasion.
@@ 3 Methodology
In this section, we introduce the key components of Theory of Mind Augmented Persuader (ToMAP). We first describe the setting of the persuasion task in Section˜3.1 and the reinforcement learning algorithm used to train the persuader in Section˜3.2. Finally, we introduce two theory of mind modules that provide critical information about the opponent’s thoughts Section˜3.3.
@@ 3.1 The Setting of Persuasion Task
The persuasion process is a multi-turn conversation between two LLM agents: a persuader, denoted as $\mathbf{X}$, and a persuadee, denoted as $\mathbf{Y}$. During the conversation, the persuader aims to convince the persuadee to support a given claim $Q$. Formally, we denote the $i$th turn in the conversation as $T_{i}$, and the conversation history till the $i$th turn as $H_{i}=T_{1\ldots i}$ [^1] Specially, we define $H_{0}$ as an empty string.. Since the two agents take turns speaking in the conversation, the next turn can be represented as:
$T_{n}\sim\begin{cases}P_{\mathbf{X}}(\cdot\mid p_{\text{debate}},Q,H_{n-1})&\quad\text{if }n\text{ is odd,}\\ P_{\mathbf{Y}}(\cdot\mid p_{\text{debate}},Q,H_{n-1})&\quad\text{if }n\text{ is even.}\end{cases}$ (1)
After each persuadee’s turn, we employ a 5-point Likert scale (Likert, 1932) to assess the persuadee’s perspectives through prompting, which can be represented as:
$s^{\prime}(\mathbf{Y},H_{n},Q)=P_{\mathbf{Y}}(\cdot\mid p_{\text{attitude}},Q,H_{n})\in\{0,1,2,3,4\}.$ (2)
In above formulas, $p_{\text{debate}}$ and $p_{\text{attitude}}$ are task-specific prompts, specified in Appendix˜A.
Since LLMs are known to exhibit inconsistency in opinions, which means a LLM might agree to both one claim and the opposite claim (Jung et al., 2022; Liang et al., 2024), we consider the persuadee’s opinion on two contradictory claims, and calculate the balanced agreement score as [^2] We have $S(\mathbf{Y},H_{n},Q)+S(\mathbf{Y},H_{n},\neg Q)=1$, which ensures the consistency of LLM opinion judgment.:
@@
where $\neg$ means the opposition of the original claim. We believe the persuadee’s self-reported attitude serves as a scalable and reliable evaluation metric for two main reasons. First, as LLMs are being deployed into practical applications, they often act as persuadees. Second, as shown in Section˜5.1, LLMs and humans demonstrate a high degree of consistency in their judgments of persuasion.
@@ 3.2 A Reinforcement Learning Formulation for Persuader Training
Reinforcement learning (RL) has demonstrated great success in recent work such as Deepseek-R1 (Guo et al., 2025), Search-r1 (Jin et al., 2025) and RM-R1 (Chen et al., 2025), where models develop complex reasoning skills. To incentivize LLMs’ persuasion ability without extensive labeled data, we also utilize RL to train the persuader through conversations and feedback from the persuadee.
As we find directly optimizing the persuader in multi-round interaction is challenging and requires much more complex design (Jin et al., 2025; Wang et al., 2025b), we propose decomposing the multi-turn conversations as single-turn activities to enhance the stability and efficiency of our training schema. Specifically, given the conversation history $H$ and target claim $Q$, we optimize the persuader to generate the next argument that maximizes the attitude shift of the persuadee. Formally, at the persuader’s turn 2k12k-1, we use
@@
to measure the attitude shift after the persuader’s turn and the persuadee’s following turn. Then, we normalize the agreement gap to the range of [1,1][-1,1] to obtain the persuasion reward: